OpenStudy (anonymous):

Felix exclaims that his quadratic with a discriminant of -1 has no real solutions. Felix then puts down his pencil and refuses to do any more work. Create an equation with a negative discriminant. Then explain to Felix, in calm and complete sentences, how to find the solutions, even though they are not real.

OpenStudy (anonymous):

I just need help with the last part on how to find the solutions.

OpenStudy (anonymous):

@zepdrix @mathmale @solidrebirth @superhelp101 @tejasvir

OpenStudy (anonymous):

haha in calm sentences

OpenStudy (anonymous):

yeah :/ it's dumb but what can you do lol

OpenStudy (anonymous):

brb.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

Woah woah woah, calm down Felix. So here's the thing... You got a discriminant of -1, \[\Large\rm b^2-4ac=-1\]

zepdrix (zepdrix):

The quadratic formula gives us solutions, both real and complex.\[\Large\rm x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]Plugging in our discriminant gives us:\[\Large\rm x=\frac{-b\pm\sqrt{-1}}{2a}\]From this point Felix, you need to recall that the imaginary unit is defined as: \(\Large\rm \mathcal i =\sqrt{-1}\) So again, we'll plug this in.\[\Large\rm x=\frac{-b\pm\mathcal i}{2a}\]We'll write this as two separate fractions,\[\Large\rm x=\frac{-b}{2a}\pm\frac{1}{2a}\mathcal i\]And here we can see that we have solutions of the form: \(\Large\rm x=a+b\mathcal i\)

zepdrix (zepdrix):

They might not be real, no, but they're solutions. So you need to sit down and cool your jets Felix.

zepdrix (zepdrix):

That's what I would tell him at least :d

zepdrix (zepdrix):

Oh they want you to actually create an equation... not leave the b and a like that. Ugh-_- thinking..

OpenStudy (anonymous):

I have part of this answered, I'll write what I wrote there, here! Give me one second it will probably help.

OpenStudy (anonymous):

To find the discriminant solutions, you need to first have the problem in standard form, like so: ax^2 + bx + c = 0 From here, you use the formula to find the discriminant, which is b^2 - 4ac. Example problem: -2x^2 + 2 - 4 = 0 Discriminant: 2^2 - 4(-2)(-4) = -28

OpenStudy (anonymous):

I'm not sure if I answered the whole question, but this is what I have so far :)

zepdrix (zepdrix):

Ok great so far. Then to actually find the solutions, you need to use the quadratic formula. And plug in your imaginary unit. The example you came up with works very well for this :)

OpenStudy (anonymous):

Okay! Can you see if I get them right, here?

zepdrix (zepdrix):

Just make sure you understand how to deal with the sqrt(-28),\[\Large\rm \sqrt{-28}=\sqrt{28}\cdot\sqrt{-1}=\sqrt{28}~\mathcal i\]

zepdrix (zepdrix):

Sure.

zepdrix (zepdrix):

In your quadratic, the middle term should have an `x`. Just making sure that was a typo.

OpenStudy (anonymous):

The solutions are -1 and 2, right?

zepdrix (zepdrix):

What? D: Nooo. We shouldn't end up with `real` solutions. Where did those numbers come from....?

OpenStudy (anonymous):

-_- i used the quadratic formula

zepdrix (zepdrix):

This is what your quadratic setup should look like. \[\Large\rm x=\frac{-2\pm\sqrt{-28}}{-4}\]

OpenStudy (anonymous):

Yeah I just realized :) Let me see if I can get it now.

OpenStudy (anonymous):

So the solutions are 1+ 2i sqr(7) / 4 and 1- 2i sqr(7) / 4

zepdrix (zepdrix):

Why a 1 for the first number?

zepdrix (zepdrix):

Wasn't that a 2 before?

OpenStudy (anonymous):

reducing!

zepdrix (zepdrix):

Reduced from where? :o Nothing else changed. Were you trying to divide it with the 4?

zepdrix (zepdrix):

So you should get here (before simplifying anything), \[\Large\rm x=\frac{-2\pm2\sqrt{7}~\mathcal i}{-4}\]From this step, yes you can divide a -2 out of everything.\[\Large\rm x=\frac{1\pm\sqrt{7}~\mathcal i}{2}\]

zepdrix (zepdrix):

Which is... close to what you have but uhhh.... yah

OpenStudy (anonymous):

so those are the "unreal" solutions, right @zepdrix ?

zepdrix (zepdrix):

Yes, those are the complex solutions. It contains both real and imaginary parts. But the solution as a whole is not all real, so we call it complex.

OpenStudy (anonymous):

Okay, awesome. Thanks for your help! :)